# Dynamic and Thermodynamic Stability of Superconducting-superfluid Stars

### v1

@@ VII Thermodynamic stability
$|v_{\text{X}}|u_{\text{X}}^{a}\delta^{2}\bm{P}_{a}^{\text{X}}-2\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\delta\bm{P}_{a}^{\text{X}}+\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\bm{P}_{a}^{\text{X}}$
@@
As an application of our results, consider a star at $T=0$ for which the entropy per electron, $s$, takes it minimum value $s=0$ throughout the star, then any perturbation for which $\delta S=0$ must have $\delta s=0$ everywhere. Similar to the single perfect fluid case as shown in [Friedman:1978wla], in this isentropic case, with $\delta S=\delta N^{\text{X}}=0$, every perturbation can be described in the Lagrangian framework. So in this case, the word “necessary” in Theorem 7 can be replaced by “necessary and sufficient”.
On the other hand, consider the spherically symmetric perturbations of a static, spherically symmetric isentropic star. For this kind of perturbations, one clearly has $\delta J^{\text{X}}=0$. Let
@@
Subjecting to the boundary condition $U^{\text{X}}=0$ at $r=0$ in the spherical coordinate leads to $U^{\text{X}}=0$ throughout the star, so there is no spherically symmetric trivial displacement. It is immediately that in the case of spherically symmetric perturbations, there is no restriction on $\mathcal{V}$ mentioned in Sec. VI, that is, $\mathcal{V}=\mathcal{C}$. Compare the Theorem 6 with the Theorem 7 (with the word “necessary” is replaced by “necessary and sufficient”), we see that in the isentropic case, for spherically symmetric perturbations of static, spherically symmetric stars, the weak thermodynamic stability is equivalent to the dynamic stability.
@@ VIII Conclusion and discussion
We have established the criterion for both dynamic and thermodynamic stability, as summarized into Theorems 4, 6, 7, and 8. To this end, we first derived the necessary and sufficient condition for the thermodynamic equilibrium, identified the degeneracy of pre-symplectic form, and constructed the phase space. This framework allowed us to derive the canonical energy, which we then established as a stability criterion by considering physically stationary solutions. As a by-product, we also derive the eigenvalue equation for the speed of sound in our superconducting-superfluid star model.
@@
There are several further directions to be investigated. First, although the vortex is allowed in our superconducting-superfluid stars, it would be beneficial to investigate the dynamic and thermodynamic stability directly from the model for dealing with the macroscopic effects of vorticity quantization [Carter:1995mj, Carter:1998rn]. Second, one could explore whether the perturbative approach employed in [Caballero:2024qtv] can be adapted to this superconducting-superfluid stellar model to yield an alternative stability criterion. Finally, it will be interesting that whether there is equivalence between the dynamic and thermodynamic stability for stars in AdS spacetimes, where the criterion for black branes in AdS spacetimes has been established in [Hollands:2012sf], or in other generalized gravitational theories.
@@ Acknowledgements.
Appendix A Speed of sound and causal behavior
The purpose of this appendix is to determine the propagation speeds and polarization directions of sound wave fronts in our superconducting-superfluid star. A convenient method is the so-called Hadamard technique (introduced long ago by Hadamard [hadamard1903leçons] and then generalized to general-relativistic elasticity theory [Rayner:1963],[bennoun1965etude]) of investigating the characteristic hypersurfaces of possible discontinuity in a partial differential system. By using such a method, Carter and Langlois succeeded in calculating the first and second sound in a relativistic two-constituent superfluid [Carter:1995if]. The method works by considering the first order case in which the algebraically related variables $n_{\text{X}}^{a}$, $s$, and $\mu_{a}^{\text{X}}$ are themselves continuous but have space-time derivatives that are weakly discontinuous across some characteristic hypersurface with tangent direction specified by some normal covector $\lambda_{a}$ say. For any component $\phi$, the discontinuities $\lceil\nabla_{a}\phi\rfloor$ in its gradient components will have to be proportional to the normal $\lambda_{a}$, i.e., we shall have $\lceil\nabla_{a}\phi\rfloor=\hat{\phi}\lambda_{a}$ for some scalar $\hat{\phi}$. Applying this to the relevant variables in the present case, we shall have
$\hat{n}_{\text{X}}^{a}\lambda_{a}$ $=0,\quad u_{\text{e}}^{a}\lambda_{a}\hat{s}=0,$
$\quad+2\sum_{\text{Z}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial n_{\text{Y}}^{2}\partial x_{\text{XZ}}^{2}}n_{\text{Z}a}n_{\text{Y}b}+\sum_{\text{Z}\neq\text{Y}}\sum_{\text{W}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial x_{\text{YZ}}^{2}\partial x_{\text{XW}}^{2}}n_{\text{W}a}n_{\text{Z}b}\bigg)\delta n_{\text{Y}}^{b}$
$\quad+\sum_{\text{Z}}\sum_{\text{Y}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial n_{\text{Z}}^{2}\partial x_{\text{XY}}^{2}}n_{\text{Y}a}n_{\text{Z}}^{b}n_{\text{Z}}^{c}+\sum_{\text{Z}\neq\text{W}}\sum_{\text{Y}\neq\text{X}}\frac{\partial^{2}\Lambda_{\text{M}}}{\partial x_{\text{WZ}}^{2}\partial x_{\text{XY}}^{2}}n_{\text{Y}a}n_{\text{Z}}^{b}n_{\text{W}}^{c}\bigg)\delta g_{bc}$
where the explicit expressions are
will be given for a suitably normalized $\lambda_{a}$ by
$\lambda_{a}=-vu_{\text{e}a}+\gamma_{a}.$ (250)
$vu_{\text{e}a}\hat{n}_{\text{X}}^{a}=\gamma_{a}\hat{n}_{\text{X}}^{a}.$ (253)
Consequently,
and
$2\gamma^{a}u_{\text{X}}^{b}\lambda_{[a}\hat{\mu}_{b]}^{\text{X}}$
The resulting eigenvalue equations for the propagation velocity $v$ is

### v2

@@ VII Thermodynamic stability
$|v_{\text{X}}|u_{\text{X}}^{a}\delta^{2}\bm{P}_{a}^{\text{X}}-2\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\delta\bm{P}_{a}^{\text{X}}$
$+\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\bm{P}_{a}^{\text{X}}$
@@
As an application of our results, consider a star at $T=0$ for which the entropy per electron, $s$, takes it minimum value $s=0$ throughout the star, then any perturbation for which $\delta S=0$ must have $\delta s=0$ everywhere. Similar to the single perfect fluid case as shown in [18], in this isentropic case, with $\delta S=\delta N^{\text{X}}=0$, every perturbation can be described in the Lagrangian framework. So in this case, the word “necessary” in Theorem 7 can be replaced by “necessary and sufficient”.
On the other hand, consider the radial perturbations of a static, spherically symmetric isentropic star. For this kind of perturbations, one clearly has $\delta J^{\text{X}}=0$. Let
@@
Subjecting to the boundary condition $U^{\text{X}}=0$ at $r=0$ in the spherical coordinate leads to $U^{\text{X}}=0$ throughout the star, so there is no spherically symmetric trivial displacement. It is immediately that in the case of radial perturbations, there is no restriction on $\mathcal{V}$, since $\mathcal{V}$ consists of all linearized solutions which is symplectically orthogonal to trivial displacements as mentioned in Sec. VI, i.e., $\mathcal{V}=\mathcal{C}$ contains all linearized solutions. Compare the Theorem 6 with the Theorem 7 (with the word “necessary” is replaced by “necessary and sufficient”), we see that in the isentropic case, for radial perturbations of static, spherically symmetric stars, the weak thermodynamic stability is equivalent to the dynamic stability.
@@ VIII Conclusion and discussion
We have established the criterion for both dynamic and thermodynamic stability, as summarized into Theorems 4, 6, 7, and 8. To this end, we first derived the necessary and sufficient condition for the thermodynamic equilibrium, identified the degeneracy of pre-symplectic form, and constructed the phase space. This framework allowed us to derive the canonical energy, which we then established as a stability criterion by considering physically stationary solutions. As a by-product, we also derive the eigenvalue equation Eq. (137) for the speed of sound in our superconducting-superfluid star model.
@@
There are several further directions to be investigated. First, although the vortex is allowed in our superconducting-superfluid stars, it would be beneficial to investigate the dynamic and thermodynamic stability directly from the model for dealing with the macroscopic effects of vorticity quantization [12, 14]. Second, in our analysis, we suppose that all neutrons and protons are in ”super” states, however, it will be more realistic to consider the case that the neutrons and/or protons carry entropy due to there are mixtures of normal and super states in these constituents. In such a case, the constituents that should be treated independently in our system will be the superfluid neutrons, the normal fluid neutrons, the superconducting protons, the (charged) normal fluid protons, and the normal fluid electrons. Finding the stability criterion in this more complex case will be helpful. Furthermore, one could also explore whether the perturbative approach employed in [6] can be adapted to this superconducting-superfluid stellar model to yield an alternative stability criterion. Finally, it will be interesting that whether there is equivalence between the dynamic and thermodynamic stability for stars in AdS spacetimes, where the criterion for black branes in AdS spacetimes has been established in [21], or in other generalized gravitational theories.